OpenStudy (anonymous):

Which of the following expressions is the 4th root of unity? a. -i b.-1 c. i d. all are correct

OpenStudy (anonymous):

@mustafa2014

OpenStudy (anonymous):

@phi

OpenStudy (phi):

1 can be written as 1(cos(0) + i sin(0) ) to find the 4th roots, divide the angle by 4 and find the 4th root of the leading 1... which is just 1 1 ( cos(0/4) + i sin(0/4) which is what we started with however there are 4 roots spaced equally around the circle 2pi/4 or pi/2 spacing so the other three roots are at angles 0+pi/2 0+ 2 pi/2, and 0 + 3 pi /2

OpenStudy (phi):

you can list all 4 roots as cos(0) + i sin(0) = 1 + i*0 = 1 cos(pi/2) + i sin(pi/2) = 0 + i*1 = i cos( pi ) + i sin(pi) = -1 + i*0 = -1 cos(3pi/2) + i sin (3pi/2) = 0 + i*-1 = -i of course, you could have just checked the choices they gave you for example: (a) -i: -i*-i * -i*-i = i^2 * i^2 = -1 * -1 = 1 works! (b) -1 * -1 * -1 * -1 = 1*1 = 1 works! ( c ) i*i * i*i = -1 * -1 = 1 works!